# MARS: A neurosymbolic approach for interpretable drug discovery

### v2

@@ A.4 Hyperparameter selection
Here, we describe hyperparameter selection. Table A2 describes the hyperparameter search space for optimization, and Table A3 describes the hyperparameters which were fixed for every model.
@@
$\lambda$ (Lambda) ratio at which the second summand is applied relative to the first summand in the reward function {5, 8, 10}
@@
hidden size size of hidden layers {64 128 256}
@@
$\gamma_{\text{baseline}}$ (gamma baseline) discount factor for the baseline {0.05, 0.5}
$\beta$ (beta) entropy regularization factor {0.025 0.05}
@@
$\gamma$ (gamma) discount factor 1
@@
A.5 PoLo metrics without associative rules
@@ A.6 Ablation study
A.7 XSwap permutations: $\text{MARS}_{P_{2H}}$ on MoA-net-10k

### v3

@@ A.6 Hyperparameter selection
Here, we describe hyperparameter selection. Table A2 describes the hyperparameter search space for optimization, and Table A3 describes the hyperparameters which were fixed for every model. Table A4 describes the best hyperparameters for the final results in Fig. 3-E.
@@
$\lambda$ (Lambda) ratio at which the second summand, or reward, is applied relative to the first summand, or reward, in the reward function {5, 8, 10}
@@
hidden size size of hidden layers {64, 128, 256}
@@
$\gamma_{\text{baseline}}$ (gamma baseline) discount factor for the baseline as implemented in MINERVA (Das et al., 2018) {0.05, 0.5}
$\beta$ (beta) entropy regularization factor as implemented in MINERVA (Das et al., 2018) {0.025, 0.05}
@@
$\gamma$ (gamma) discount factor as implemented in REINFORCE (Williams, 1992) 1
@@
Table A4: Best hyperparameters from Table A2 for the experiments in Fig. 3-E.
Hyperparameter $\text{MARS}_{P_{2H}}$ $\text{MARS}_{\text{naive}}$ PoLo
$\lambda$ 10 5 5
$\alpha$ 0.001 0.001 -
learning rate 0.0001 0.0001 0.0001
hidden size 256 256 64
batch size 128 256 256
rollouts 100 100 50
$\gamma_{\text{baseline}}$ 0.05 0.5 0.5
$\beta$ 0.025 0.05 0.05
A.7 Additional performance metrics
Within Fig. 7, we report Hits@1 and Hits@3 as in Fig. 3:
Figure 7: Hits@1 and Hits@3 for $\text{MARS}_{P_{2H}}$ compared to PoLo and $\text{MARS}_{\text{naive}}$ upon several variants of MoA-net. Each bar is the average and standard deviation across five independent training and testing iterations. From left to right: Little change between initial metrics upon MoA-net (A) in comparison to the standard MoA-net-permuted metrics (B) provides evidence that predictions are influenced by degree bias, resulting in a reasoning shortcut. Thereafter, inverse edges were removed to prohibit the reasoning shortcut, hindering performance (C). Performance was restored upon MoA-net-10k with the KG trimming step (D), with $\text{MARS}_{P_{2H}}$ showing the best standard and pruned metrics. Finally, $\text{MARS}_{P_{2H}}$ maintains high pruned metrics even when inverse edges (and reasoning shortcuts) are re-introduced (E).
A.8 PoLo metrics without associative rules
@@ A.9 Ablation study
A.10 XSwap permutations: $\text{MARS}_{P_{2H}}$ on MoA-net-10k